solve x to 4 decimal places 5^x=4^(x+1)
Answer by rfadrogane(214)

solve x to 4 decimal places 5^x=4^(x+1)
sol'n:
5^x = 4^(x+1), take ln on both side
x ln 5 = (x+1)ln 4
x ln 5 = x ln 4 + ln 4
x (ln 5 - ln 4)= ln 4
thus;
x = ln 4/(ln 5 - ln 4)
x = 6.2126 ----answer